OpenStudy (anonymous):

One number is 2 less than two-thirds the other. Their sum is 103. Find the two numbers

hero (hero):

n = 1st number m = 2nd number n = 2m/3 - 2 n + m = 103 (2m/3 - 2) + m = 103 2m/3 + m - 2 = 103 2m/3 + m = 103 + 2 2m/3 + m = 105 2m/3 + 3m/3 = 105 5m/3 = 105 5m = 315 m = 315/5 m = 63 n = 40

OpenStudy (anonymous):

thank you